OpenStudy (vval25):

verify that y=e^-x is a solution to the differential equation 3y'+4y=e^-x

OpenStudy (phi):

do you know how to take the derivative (with respect to x) of e^-x ?

OpenStudy (vval25):

Yes, (-1/e^x)

OpenStudy (elonasushchik):

@phi

OpenStudy (phi):

I would leave it -e^-x (not that it matters too much) replace y' with -e^-x replace y with e^-x and check that the left side matches the right side in your equaiotn

OpenStudy (vval25):

yes, I did that, then what

OpenStudy (freckles):

you are trying to see if both sides are the same when you plug in your y' and y with -e^(-x) and e^(-x) respectively

OpenStudy (freckles):

if both sides are the same then y=e^(-x) is a solution if both sides aren't the same then y=e^(-x) is not a solution

OpenStudy (vval25):

I get -3/e^x +4e^-x = e^-x

OpenStudy (vval25):

I got the answer, thank you everyone